OpenStudy (anonymous):

help please?!? 8th term: 22, 19.5, 17, ...

OpenStudy (anonymous):

common difference is -2.5 so, 8th term= 22-2.5x7=22-17.5=4.5

OpenStudy (agent0smith):

Do you know how to find the common difference between terms? It's the amount that is added to a term to get the next term. you can use the formula for an arithmetic sequence: \[\Large a _{n} = a _{1} + (n – 1)d\] a1 is the first term in the sequence, d is the common difference. You want the 8th term, so n = 8.